OpenStudy (anonymous):

CHECK MY ANSWER PLEASE! 42.5 grams of an unknown substance is heated to 105.0 degrees Celsius and then placed into a calorimeter containing 110.0 grams of water at 24.2 degrees Celsius. If the final temperature reached in the calorimeter is 32.4 degrees Celsius, what is the specific heat of the unknown substance? Show or explain the work needed to solve this problem, and remember that the specific heat capacity of water is 4.18 J/(°C x g).

OpenStudy (anonymous):

4.18*103.5*(32.5-24.5)=c(31.5)(102.4-32.5) 432.63(5)=C(31.5)(69.9) 2163.15=C2201.85 C=.982

OpenStudy (aaronq):

your numbers don't match the ones given, where did the 103.5 and 31.5 come from?